OpenStudy (anonymous):

Solve the system by elimination method. 0.5x=-2.0+y , -4/3x+y=-29/3

OpenStudy (anonymous):

Let's organize the system. {0.5x-y=-2 {-4/3x+y=-29/3 Let's multiply the second equation by 3 and the first by 2. {x-2y=-4 {-4x+3y=-29 Now we eliminate x by multiplying the first equation by 4. {4x-8y=-16 {-4x+3y=-29 We add them. -5y=-45. y=9. so that means x-2(9)=-4. 14-18=-4. So x=14. The answer is (14,9).

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

Youre welcome. Thanks to you, I got a new title!(I reached 50 in problem solving)

OpenStudy (anonymous):

yay!